A 91.0-kg mail bag hangs by a vertical rope 3.9 m long. A postal worker then displaces the bag to a position 2.9 m sideways from its original position, always keeping the rope taut.

Part A

What horizontal force is necessary to hold the bag in the new position?

Express your answer with the appropriate units.

Part B

As the bag is moved to this position, how much work is done by the rope?

Express your answer with the appropriate units.

Part C

As the bag is moved to this position, how much work is done by the worker?

Express your answer with the appropriate units.A 91.0-kg mail bag hangs by a vertical rope 3.9 m long. A postal worker then displaces the bag to a position 2.9 m sideways fHi, please make sure your answers are correct.
Answer #1

sino=d/e = 2 2.9 3.9 = 0.7435 a) O = sin(0.7435) 0 - 48.039 F = mg tano = (91)(9.8) tan (48.034) F = 989.898N W = Tdcosgol Te